OpenStudy (anonymous):

please help!!!!! fast. perform the indicated operation. 6 and 1/3rd plus 8 and 3/4ths

OpenStudy (anonymous):

46/3

OpenStudy (anonymous):



OpenStudy (♪chibiterasu):

We have to get the denominators to be equal, so we have to multiply them to the greatest common factor. The GCF of 3 and 4 is 12, so we have to multiply both fractions by 4 and 3: \[\frac { 3 }{ 4 } • 3 = \frac { 9 }{ 12 }\] \[\frac { 1 }{ 3 } • 4 = \frac { 4 }{ 12}\] Now: \[6 \frac { 4 }{ 12 } + 8 \frac{ 9 }{ 12 } = 12 \frac { 13 }{ 12 }\]

OpenStudy (♪chibiterasu):

13/12 = 1 1/12, so the answer is \[15 \frac{ 1 }{ 12 }\]

OpenStudy (♪chibiterasu):

I meant 14 13/12, not 12/ 13/12.